OpenStudy (anonymous):

how do i find the equation of the parabola passing through the points (1,0), (3,0) and (2,-1)?

OpenStudy (precal):

y=1x^2-4x+3

OpenStudy (amistre64):

notice that your roots are at 1 and 3 so the vertex is centered between them at x=2 for starters

OpenStudy (precal):

Also, it is concave up so therefore it is a Positive x^2

OpenStudy (anonymous):

what is the procedure you did to find it precal?

OpenStudy (amistre64):

multiply your roots and scale it to fit

OpenStudy (precal):

I cheated. I just used my TI-84 + and did a regression

OpenStudy (amistre64):

(x-1)(x-3) = x^2-4x +3 see how far off f(2) is from -1

OpenStudy (precal):

That is your axis of symmetry

OpenStudy (amistre64):

f(2) = 2^2 -4(2) + 3 = 4 -8 + 3 = -1 its not off so its just right